# Predicting the Performance of Black-box LLMs through Self-Queries

### v2

@@ 4 Experiments
We now evaluate the utility of these extracted representations in three main applications: (1) predicting the performance of various open- and closed-source LLMs on a variety of text classification and generation tasks, (2) detecting whether a LLM has been influenced by an adversary, and (3) distinguishing between different LLM architectures. We refer to our approach as QueRE (Question Representation Elicitation). Code to replicate our experiments can be found at https://github.com/dsam99/QueRE.
@@
In our experiments, we compare against a variety of different baselines; two of which are strong baselines that assume access to more information than our approach. These are RepE (Zou et al., 2023a), which extracts the hidden state of the LLM at the last token position in its representation reading, and Full Logits, which uses the distribution over the LLM’s entire vocabulary. Both of these cannot be applied to black-box language models and should be seen as strong comparisons that assume more information than our approach. For instance, information from the full logits over the complete vocabulary has been shown to reveal hyperparameter information of the LLM (Finlayson et al., 2024). For this baseline for black-box models, we best approximate this with a sparse vector of the top-k probabilities (if that is provided by the API). We also compare against a version of semantic uncertainty (Kuhn et al., 2023) on the closed-ended QA tasks, which looks to extract a more accurate quantification of uncertainty by grouping together semantically similar tokens for each potential answer. This baseline does not straightforwardly apply to open-ended QA tasks, and we only present results for semantic uncertainty on the open-source models as we do not have access to all of the GPT model’s token probabilities.
We also compare against pre-conf and post-conf scores, which are a univariate feature that corresponds to the probability of the “yes” token under the language model to a prompt about the model’s confidence either before (pre-) or after (post-) seeing the greedy (temperature 0) sampled response. This is the same as the naive approach in directly extracting confidence scores from LLMs (Xiong et al., 2023). We also compare against using the normalized probability distribution over the potential answer questions (Answer Probs), which is similar to what is proposed in prior work that focuses on in-context learning (Abbas et al., 2024). These are individual components of our representations, so these comparisons illuminate how much of an increase in performance we obtain by adding additional elicitation queries and concatenating them together.
@@
We compare our approach to the baselines on a variety of open-ended QA tasks (NQ (Kwiatkowski et al., 2019), SQuAD (Rajpurkar et al., 2016)) and closed-ended QA tasks (BoolQ (Clark et al., 2019), WinoGrande (Sakaguchi et al., 2021)), including detecting hallucinations (HaluEval (Li et al., 2023)) and toxic comments (DHate (Vidgen et al., 2021)), and commonsense reasoning (CS QA (Talmor et al., 2019)). In our experiments, we evaluate the performance of LLaMA3 (3B, 8B, and 70B) (Dubey et al., 2024) and OpenAI’s GPT-3.5 and GPT-4o-mini models (Achiam et al., 2023). In all of the text generation tasks, we sample greedily from the LLM for its answer. On the NQ dataset, we prepend prompts with two held-out training examples to have the LLMs better match the answer format. Additional experimental details can be found in Section D.5.
Table 1: Accuracy in detecting if GPT models have been adversarially influenced by a system prompt on QA and code generation tasks. For QA, we use linear models on top of each method to distinguish between 3 copies of clean and 3 copies of adversarially influenced LLMs via a system prompt. For code generation, we use linear models to detect between a clean and an adversarial LLM that will secretly introduce bugs according to system prompts (see Table 1 for more details.)
Dataset Model Pre-conf Full Logits QueRE
BoolQ GPT-3.5-turbo 0.5396 0.8483 0.8668
GPT-4o-mini 0.5725 0.9033 0.9547
CodeContests GPT-3.5-turbo 0.5061 0.9455 0.9909
GPT-4o-mini 0.4454 0.8848 1.0000
4.1 Predicting Model Performance
Our first evaluation is to predict the performance of the LLM on QA tasks. Accurately predicting model performance has many benefits, such as enabling better resource allocation by identifying challenging tasks and mitigating potential failures in high-stakes environments where incorrect predictions have significant consequences. For open-ended QA tasks, we measure if the model has produced a correct answer under other metrics. For instance, on SQuAD (Rajpurkar et al., 2016), we measure if the model has produced the exact match, and on Natural Questions (NQ) (Kwiatkowski et al., 2019), we measure if the LLM has outputted one of the valid answers to the question. For closed-ended QA tasks, we simply predict the 0-1 error.
Overall, we observe that QueRE strongly outperforms all other methods (including white-box approaches) on open-ended QA tasks (Figure 2) and is most often the best performing black-box method on closed-ended tasks (Figure 3). We compare with other uncertainty quantification approaches from (Xiong et al., 2023) in Section A.2, similarly finding that QueRE strongly outperforms using a variety of different uncertainty quantification techniques. We defer results on the remaining models to Appendix A.1, where similar trends hold. These results suggest that our approach results in useful representations, even when compared to white-box baselines, where useful information about model behavior are extracted through our use of follow-up queries.
@@ 4.1 Predicting Model Performance
4.2 Detecting Adversarial/Harmful LLMs
@@
Next, we demonstrate QueRE can reliably distinguish between a clean version of the LLM and one that has been influenced by an adversary, which is crucial for maintaining robust and trustworthy AI systems. We provide an experiment on both a QA task and a code generation task, where we add an adversarial system prompt to instruct GPT models to answer questions or generate code incorrectly. For the QA task, we use multiple copies of the LLM with 3 different adversarial and 3 different helpful system prompts, and train linear probes on top of the extracted features to distinguish between helpful and harmful system prompts (see Section D.1 for more details and for the specific prompts). For the code generation task, we use multiple copies of the LLM with different adversarial system prompts to generate incorrect code on the Code Contests dataset (Li et al., 2022). These experimental settings are similar to the work of MacDiarmid et al. (2024), where they could reliably detect the presence of sleeper agents, except in our setting we assume only black-box access to the model.
@@
Our results show that QueRE (with a simple linear probe) can reliably detect when this a model has been influenced by an adversarial system prompt in both QA and code generation settings (Table 1), outperforming or matching all other methods on all datasets. These results also suggest that QueRE is also robust to the specific framing of the adversarial prompt, given that it can distinguish between multiple versions of harmful versus helpful system prompts in both QA and code generation tasks.
4.3 Distinguishing Between Model Architectures
@@
Finally, we consider the setting of distinguishing between different model architectures in a black-box setting, purely via analyzing their outputs. This has a practical application; when using models given through an API, our approach can be used to reliably detect whether a cheaper, smaller model is being falsely provided through an API, as is the focus of prior work (Chen et al., 2023). This problem has also been studied by concurrent work (Gao et al., 2024) in the setting of hypothesis testing. We provide an experiment where the goal is to classify which LLM from which each extracted representation was generated.
@@
We demonstrate that QueRE can be used to reliably distinguish between different LLM architectures and sizes. Figure 4. We observe that linear predictors using QueRE can often almost perfectly classify between LLMs of different sizes, while most other approaches often struggle with this task. This suggests that the distributions learned by different LLMs behave in distinct ways, even when the same architecture and training objectives are used and the only difference is the model size. Notably, this suggests that different model scales cannot be differentiated simply through naive confidence scores.
@@
Table 2: Transferrability of representations to OOD settings, where we train linear classifiers to predict model performance on one QA task and transfer to another target QA task. The dataset transfer is run for LLaMA3-70B. The model transfer is run on SQuAD, and we do not report results for RepE as model activations are of different sizes. We report results in terms of AUROC.
@@
Transfer Full Logits RepE Pre-conf Post-conf Answer Probs QueRE
Squad $\rightarrow$ NQ 0.5716 0.4896 0.5563 0.7976 0.8328 0.8964
NQ $\rightarrow$ Squad 0.5283 0.4967 0.5099 0.7818 0.7532 0.7934
LLaMA3-3B $\rightarrow$ 8B 0.5477 0.5145 0.7928 0.6061 0.8409
LLaMA3-8B $\rightarrow$ 70B 0.4880 0.5099 0.7818 0.7532 0.8295
Figure 5: Expected Calibration Error (ECE) for QueRE and Answer Probs on HaluEval (Left) and SQuAD (Right); lower values are better. We observe models trained on QueRE are more calibrated.
@@
Table 3: Generalization bounds on accuracy in predicting model performance on QA tasks. We bold the best (highest-valued) lower bound on accuracy. We use $\delta=0.01$.
@@
Dataset LLM Full Logits RepE Answer Probs QueRE
SQuAD LLaMA3-8B 0.5979 0.5728 0.5126 0.8088
LLaMA3-70B 0.4996 0.4496 0.5881 0.7558
BoolQ LLaMA3-8B 0.6011 0.5521 0.4608 0.6706
LLaMA3-70B 0.5616 0.5590 0.3602 0.5232
@@
Figure 6: AUROC as we vary the number of random samples $k$ used to approximate LLM probabilities with GPT-3.5 on HaluEval (left) and DHate (right) over 5 random seeds. We observe that there is not a significant dropoff in performance when using approximations due to sampling.
Figure 7: AUROC on predicting model performance with our black-box representations on BoolQ for LLaMA3-8B (left) and LLaMA3-70B (right). The shaded area represents the standard error.
@@ 4.4 Additional Results
Transferability of Extracted Representationss
We also provide experiments that demonstrate the generality and transferrability of classifiers trained on representations extracted via QueRE to OOD settings. We present the comparison of QueRE against the other baselines as we transfer from one QA dataset to another (using the LLaMA3-70B model). We also provide an experiment as we transfer from representations extracted from one LLaMA3 model size to another. Across all tasks, QueRE shows the best transferring performance (Table 2). Thus, ours is the best approach for tackling these OOD settings without any access to labeled data from the target task.
Calibration
While we have previously reported the AUROC of our predictors, we are also interested in the calibration of our models (e.g., accuracy at a given confidence threshold). This is particularly useful for high-stakes settings, when we may only want to defer prediction to a LLM when we are confident in its performance. We observe that predictors defined by QueRE generally have much lower ECE compared to those defined by using answer probabilities. We defer results on other datasets to Appendix A.6. Our approach shows promise in constructing well-calibrated and performant predictors of LLM performance, which are important for the application of LLMs in high-stakes settings (Weissler et al., 2021; Thirunavukarasu et al., 2023).
Generalization Bounds
Another added benefit of our approach is that it yields low-dimensional representations, which can be used with simple models, to achieve strong predictors of performance with tight generalization bounds. Bounds for linear models that use features from a pretrained model have been explored in practice (McNamara and Balcan, 2017), although not for LLMs. Another key difference is that, while we similarly extract a representation from the model, previous approaches use a penultimate layer rather than the ability of a LLM to generated features in response to language queries. We use the following PAC-Bayes generalization bound for linear models (see Section A.5 for more details). We observe that linear predictors trained our representations have stronger guarantees on accuracy, when compared to baselines (Table 3 and Appendix A.5). A limitation of these results is that they require an assumption that the representations extracted by a LLM are independent of the downstream task data; this assumption is verifiable via works in data contamination (Oren et al., 2023) or is valid on datasets released after LLM training (e.g., HaluEval for GPT-3.5).
@@ 4.5 Ablations
Sampling from the Black-Box LLM Achieves Comparable Performance
As previously mentioned, we often do not have access to top-$k$ probabilities through the closed-source API. While we have provided asymptotic guarantees (in terms of both $n$ and $k$) on the estimator learned via logistic regression, we are also interested in the setting where we have a finite number of samples $k$. Therefore, we run an experiment where instead of using the actual ground-truth probability, we approximate this via an average of $k$ samples from the distribution of the LLM. We report results using approximations via sampling from the distribution specified by GPT-3.5’s top-$k$ log probs (Figure 6). We do not observe a significant drop (less than 2 points in AUROC) in performance when using sampling, which implies that our method can be used in settings with closed-source LLMs that do not give top-$k$ probability access.
More Elicitation Questions Leads to Better Performance
We study how much the number of elicitation questions directly impacts how much information is extracted in QueRE. We randomly subsample the number of elicitation questions and report how much the performance of our approach varies when only using this subset of questions. We observe the overall trend that our predictive performance increases as we increase the number of elicitation prompts (Figure 7), with the rate of increase slowly diminishing with more prompts. We defer results on other datasets to Appendix A.10, where we observe similar results. Overall, this demonstrates that we can achieve even stronger performance with our method with more elicitation questions, even with those that are LLM-generated.
Figure 8: Comparison of using varying amounts of prompts of random sequences of natural language or elicitation questions in QueRE. The results for the LLaMA3-70B model are presented from left-to-right as: SQuAD, NQ, and HaluEval. Results are averaged over 20 random seeds.
Elicitation Questions Versus Random Sequences of Language
We also analyze the impact of the importance of the particular choice of our elicitation questions (i.e., generated via GPT-4 in a certain way) by running an ablation study where we feed unrelated sequences of coherent natural language as inputs to the model. This new comparison evaluates how much random sequences of natural language influence the distribution from the LLM and studies how useful this extracted information is for downstream tasks (see Appendix D.3 and Appendix D.4 for more details). We present some of these experiments in Figure 8, and defer the remaining results to Appendix A.4. While using these unrelated sequences in QueRE often leads to worse performance than using meaningful elicitation questions, the observation that responses to unrelated sequences give useful and generalizable information about a model’s decision and can even sometimes outperform interpretable questions is an interesting result. This suggests that additional follow-up prompts can be easily generated, as they do not need to be in the form of meaningful questions to reveal information about model behavior, still providing notions of diversity. We also study the impacts of different types of elicitation questions in Section A.3, again finding that human-interpretable forms of diversity do not necessarily lead to more diverse features from the LLM.

### v4

@@ 4 Experiments
We now evaluate our method in three main applications: (1) predicting the performance of various open- and closed-source LLMs on a variety of text classification and generation tasks, (2) detecting whether a LLM has been influenced by an adversary, and (3) distinguishing between different LLM architectures. We refer to our approach as QueRE (Follow-up Question Representation Elicitation).
@@
In our experiments, we compare against a variety of different baselines. Our first two baselines are white-box methods, which assume more information than QueRE. These include RepE (zou2023representation), which extracts the hidden state of the LLM at the last token position, and Full Logits, which uses the distribution over the LLM’s entire vocabulary. Neither of these can be applied to black-box language models and should be seen as strong baseline comparisons. For instance, information from the full logits over the complete vocabulary has been shown to reveal proprietary information from LLMs (finlayson2024logits). To approximate Full Logits for black-box LLMs, we approximate this with a sparse vector of top-k probabilities provided by the API.
For black-box baselines on open-ended QA tasks, we compare against Self-Consistency (wei2024measuring), where we sample 10 times from the language model to define a probability distribution over potential answers. For closed-ended QA tasks, we can directly use the probability distribution over the potential answer questions (Answer Probs), as is done in prior work (abbas2024enhancing). We also compare with Semantic Entropy (kuhn2023semantic) on open-ended tasks, which aims to extract a more accurate quantification of uncertainty by grouping semantically similar answers. Finally, on all tasks, we also compare against pre-conf and post-conf scores, which are a univariate feature that corresponds to the probability of the “yes" token under the language model to a prompt about the model’s confidence either before (pre-) or after (post-) seeing the greedy (temperature 0) sampled response. This is the same as the naive approach in directly extracting confidence scores from LLMs (xiong2023can). Pre- and post-conf (and Answer Probs on closed-source tasks) are components of our representations on closed-source tasks, so this comparison illuminates how much of our performance is gained by our follow-up queries.
@@
We evaluate predicting the behavior of LLMs on a variety of benchmarks. We consider the open-ended QA benchmarks NQ (47761) and SQuAD (rajpurkar2016squad)), as well as the closed-ended QA datasets of BoolQ (clark2019boolq), WinoGrande (sakaguchi2021winogrande), HaluEval (HaluEval), DHate (vidgen2021learning), and CS QA (talmor2019commonsenseqa)). These datasets encompass commonsense reasoning, hallucination detection, factual recall, and toxicity classification. Finally, we also evaluate on math (GSM8K (cobbe2021gsm8k)) and code (Code Contests (doi:10.1126/science.abq1158)) benchmarks to evaluate if our approach is predictive of tasks that require reasoning. In our experiments, we evaluate the performance of LLaMA3 (3B, 8B, and 70B) (dubey2024llama) and OpenAI’s GPT-3.5 and GPT-4o-mini models (achiam2023gpt). In all of the text generation tasks, we sample greedily from the LLM for its answer. Additional experimental details can be found in Section˜D.5.
4.1 Predicting Model Correctness on QA and Reasoning Tasks
Our first evaluation focuses on predicting instance-level LLM performance on QA and reasoning benchmarks, according to each benchmark’s respective metric. For example, on SQuAD (rajpurkar2016squad), correctness is defined by exact match, while for reasoning benchmarks such as math and code, correctness is determined using GPT-4o as an LLM judge.
We find that QueRE consistently outperforms other methods (including white-box approaches) on open-ended QA tasks (Figure˜2) and is most often the best-performing black-box method on closed-ended QA tasks (Figure˜3). While we do not claim that QueRE captures semantic notions of reasoning, it nevertheless proves highly predictive of performance on reasoning tasks (e.g., coding and math benchmarks), while other approaches fail. Full results across all models are provided in Appendix A.1, where similar trends hold. We also compare QueRE to other uncertainty quantification approaches from (xiong2023can) in Section˜A.3, similarly finding that QueRE outperforms these techniques as well. Overall, our approach in using follow-up queries leads to predictive features for a wide variety of tasks, often rivalling or exceeding white-box baselines.
Table 2: Accuracy in detecting if GPT models have been adversarially influenced by a system prompt on QA and code generation tasks. On BoolQ, the LLMs has been influenced to answer questions incorrectly. On CodeContests, the LLM has been instucted to secretly introduce bugs into generated code. QueRE accurately detects adversarially influenced LLMs.
Dataset LLM Pre-conf Post-Conf Logits Sem. Entropy QueRE
BoolQ GPT-3.5-turbo 0.5396 0.7483 0.8483 0.5928 0.8668
GPT-4o-mini 0.5725 0.6111 0.9033 0.6134 0.9258
CodeContests GPT-3.5-turbo 0.5061 0.6515 0.9455 0.5287 0.9909
GPT-4o-mini 0.5546 0.5333 0.8848 0.6518 1.0000
@@ 4.1 Predicting Model Correctness on QA and Reasoning Tasks
4.2 Detecting Adversarially Influenced LLMs
@@
Next, we demonstrate QueRE can reliably distinguish between a clean version of the LLM and one that has been influenced by an adversary, which is crucial for maintaining robust and trustworthy AI systems. We provide an experiment on both a QA task and a code generation task, where we add an adversarial system prompt to instruct GPT models to answer questions or generate code incorrectly. For the QA task, we use multiple copies of the LLM with 3 different adversarial and 3 different helpful system prompts, and train linear probes on top of the extracted features to distinguish between helpful and harmful system prompts (see Section˜D.1 for more details and for the specific prompts). For the code generation task, we use multiple copies of the LLM with different adversarial system prompts to generate incorrect code. These experimental settings are similar to the work of macdiarmid2024sleeperagentprobes, where they could reliably detect the presence of sleeper agents in the white-box setting. In our setting, we only assume black-box access to the model.
@@
Our results show that QueRE (with a simple linear probe) can reliably detect when a model has been influenced by an adversarial system prompt in both QA and code generation settings (Table˜2), outperforming other methods on all datasets. These results also suggest that QueRE is also robust to the specific framing of the adversarial prompt, given that it can distinguish between multiple versions of harmful versus helpful system prompts in both QA and code generation tasks.
4.3 Distinguishing Between Black-box LLMs
@@
Finally, we consider the setting of distinguishing between different LLMs in a black-box setting, purely via analyzing their outputs. This has a practical application; when using models given through an API, our approach can be used to reliably detect whether a cheaper, smaller model is being falsely provided through an API. This problem has also been studied by concurrent work (gao2024model) in the setting of hypothesis testing. We provide an experiment where the goal is to classify which LLM from which each extracted representation was generated.
@@
We demonstrate that QueRE can be used to reliably distinguish between different LLM architectures and sizes (Figure˜4 and in Appendix section˜A.7). We observe that linear predictors using QueRE can often almost perfectly classify between LLMs of different sizes, while other black-box approaches do not perform as well. This suggests that the distributions learned by different LLMs behave in distinct ways, even within the same family, and the only difference is the model size. Notably, this suggests that different model scales cannot be differentiated simply through naive confidence scores.
@@
Table 3: Transferability of representations to OOD settings, where we either train linear classifiers to predict model performance on one QA task and (1) transfer to another target QA task or (2) transfer to a different QA dataset. The dataset transfer is run for LLaMA3-70B. The model transfer is run on SQuAD, and we do not report results for RepE as model activations are of different sizes. QueRE performs the best when transferred across models or datasets.
@@
Transfer Full Logits RepE Pre-conf Post-conf Self-Consis. Sem. Entropy QueRE
Squad $\rightarrow$ NQ 0.5716 0.4896 0.5563 0.7976 0.8328 0.6661 0.8964
NQ $\rightarrow$ Squad 0.5283 0.4967 0.5099 0.7818 0.7532 0.5013 0.7934
3B $\rightarrow$ 8B 0.5477 0.5145 0.7928 0.4635 0.6328 0.8409
8B $\rightarrow$ 70B 0.4880 0.5099 0.7818 0.5280 0.6658 0.8295
@@
Table 4: Generalization bounds in predicting model performance on QA tasks. We bold the best (highest-valued) lower bound on accuracy. We use $\delta=0.01$.
@@
Dataset LLM Full Logits RepE Self-Consis. Sem. Entropy QueRE
NQ LLaMA3-8B 0.4622 0.4525 0.3868 0.4534 0.7409
LLaMA3-70B 0.4752 0.4684 0.3036 0.4379 0.6495
SQuAD LLaMA3-8B 0.5979 0.5728 0.4544 0.3048 0.8088
LLaMA3-70B 0.4996 0.4496 0.2929 0.2931 0.7558
@@
Figure 5: Left: AUROC as we vary the number of random samples $k$ used to approximate LLM probabilities with GPT-3.5 on HaluEval over 5 random seeds. We observe that there is not a significant dropoff in performance when using approximations due to sampling. Right: AUROC on predicting LLaMA3-70B performance on BoolQ with QueRE as we increase the number of follow-up questions. The shaded area represents the standard error.
@@ 4.4 Additional Results
We present additional results on the generality of our approach through its ability to transfer across different datasets and models, as well as yield tight generalization bounds. We defer further results on the improved calibration of predictors learned via QueRE to Appendix A.4.
QueRE transfers across datasets and models.
We also provide experiments that demonstrate the generality and transferrability of classifiers trained on representations extracted via QueRE to OOD settings. We compare QueRE to other baselines as we (1) transfer the learned predictors from one QA dataset to another, or (2) transfer from one LLaMA3 model size to another. Across all tasks, QueRE shows the best transferring performance (Table˜3). Thus, this suggests QueRE performs the best in OOD settings without any access to labeled data from the target task.
QueRE yields tighter generalization bounds.
Another added benefit of our approach is that it yields low-dimensional representations, which can be used with simple models, to achieve strong predictors of performance with tight generalization bounds. We use the following PAC-Bayes generalization bound for linear models (see Section˜A.8 for more details). We observe that linear predictors trained our representations have stronger guarantees on accuracy, when compared to baselines (Table˜4 and Appendix A.8). A limitation of these results is that they require an assumption that the representations extracted by a LLM are independent of the downstream task data; this assumption is verifiable via works in data contamination (oren2023proving) or is valid on datasets released after LLM training (e.g., HaluEval for GPT-3.5).
@@ 4.5 Ablations
Sampling-based approximations achieve comparable performance.
As previously mentioned, we often do not have access to top-$k$ probabilities through the closed-source API. While we have provided asymptotic guarantees (in terms of both $n$ and $k$) on the estimator learned via logistic regression, we are also interested in the setting where we have a finite number of samples $k$. Therefore, we run an experiment where instead of using the actual ground-truth probability, we approximate this via an average of $k$ samples from the distribution of the LLM. We report results using approximations via sampling from the distribution specified by GPT-3.5’s top-$k$ log probs (Figure˜5 - Left). We do not observe a significant drop (less than 2 points in AUROC) in performance when using sampling, which implies that our method can be used with APIs that do not provide top-$k$ probabilities.
More follow-up questions lead to better performance.
We study how much the number of elicitation questions directly impacts how much information is extracted in QueRE. We randomly subsample the number of elicitation questions and report how much the performance of our approach varies when only using this subset of questions. We observe the overall trend that our predictive performance increases as we increase the number of elicitation prompts (Figure˜5 - Right), with the rate of increase slowly diminishing with more prompts. We defer results on other datasets to Appendix A.12, where we observe similar results. Overall, this demonstrates that we can achieve even stronger performance with our method by scaling up the number of follow-up questions. As previously mentioned, this only comes with a slight increase in computational complexity, as these follow-up questions can all be handled in parallel.
We defer further ablations on using MLPs instead of linear models in Appendix A.11 and on the type of follow-up questions used in QueRE to Appendix A.5.